# Central Answer Modeling for an Embodied Multi-LLM System

### v4

@@ Abstract
Embodied Question Answering (EQA) is an important problem, which involves an agent exploring the environment to answer user queries. In the existing literature, EQA has exclusively been studied in single-agent scenarios, where exploration can be time-consuming and costly. In this work, we consider EQA in a multi-agent framework involving multiple large language models (LLM) based agents independently answering queries about a household environment. To generate one answer for each query, we use the individual responses to train a Central Answer Model (CAM) that aggregates responses for a robust answer. While prior Question Answering (QA) work has used a central module based on answers from multiple LLM-based experts, we specifically look at applying this framework to embodied LLM-based agents that must physically explore the environment first to become experts on their given environment to answer questions. Our work is the first to utilize a central answer model framework with embodied agents that must rely on exploring an unknown environment. We set up a variation of EQA where instead of the agents exploring the environment after the question is asked, the agents first explore the environment for a set amount of time and then answer a set of queries. Using CAM, we observe a $46\%$ higher EQA accuracy when compared against aggregation methods for ensemble LLM, such as voting schemes and debates. CAM does not require any form of agent communication, alleviating it from the associated costs. We ablate CAM with various nonlinear (neural network, random forest, decision tree, XGBoost) and linear (logistic regression classifier, SVM) algorithms. We experiment in various topological graph environments and examine the case where one of the agents is malicious and purposes contributes responses it believes to be wrong. Finally, we present a feature importance analysis for CAM via permutation feature importance (PFI). We provide and code and query dataset for further research.
Introduction
Embodied Question Answering (EQA) is an important task where an agent explores an environment, such as a household, to answer questions based on its observations. It has gained considerable interest in the past few years due to its applications for in-home robots and personalized assistants (Das et al. 2018; Gordon et al. 2018; Duan et al. 2022a; Sima et al. 2023; Luo et al. 2023; Dorbala et al. 2024a; Zhu et al. 2023). For a seamless interface with humans, robotic agents must be equipped with natural language understanding (Xu et al. 2021; Sun et al. 2024) and possess common-sense reasoning of the world. One form of common sense is having a relational understanding of household items and their usual place (i.e., a mug is highly likely to be in the kitchen or living room). Foundation models, specifically Large Language Models (LLMs) such as GPT (Achiam et al. 2023) and LLama (Touvron et al. 2023), have shown remarkable prowess in both natural language understanding and common sense reasoning. Recently, prior art has used these foundation models in EQA for query generation (Dorbala et al. 2024a) and answering (Majumdar et al. 2024).
@@ Introduction
In existing works, only a single agent explores the environment with a limited field of view, inducing low coverage and high exploration costs. In contrast, for an LLM-based embodied agent to succeed in EQA, efficient exploration of the environment is needed as the questions will center around the state of the environment (Das et al. 2018; Yu et al. 2019; Zhu et al. 2023; Majumdar et al. 2024). Prior work has studied the zero-shot capabilities of LLM-based agents to explore an unknown household environment (Dorbala, Mullen Jr, and Manocha 2023; Dorbala et al. 2024b; Majumdar et al. 2024).
@@
The rise of Embodied Artificial Intelligence (EAI) (Duan et al. 2022b) brings up the possibility of co-habited spaces with multiple agents and humans working in conjunction (Puig et al. 2023). In this context, we consider a similar co-inhabited setup, and propose using a team comprising of multiple LLM-based agents or a Multi-LLM system for EQA. A Multi-LLM system allows for simultaneous exploration of the environment in the same amount of time.
@@
A key challenge in a multi-agent setup is with conflicting responses when agents are asked the same query. Since LLMs are blackboxes capable of generating free-form natural language, their responses to the same query may significantly vary (Gal and Grosz 2022; Wang et al. 2024). In an embodied setting, such responses are further skewed by partial observations due to limited exploration. Prior work in ensemble LLM methods tackle this issue using weighted sum or voting schemes (Li et al. 2024; Wei et al. 2022; Wang et al. 2022; Yang et al. 2024) and consensus-reaching debates between the LLM agents (Chen et al. 2023a; Du et al. 2023; Xiong et al. 2023; Guo et al. 2024a; Chan et al. 2023; Li, Patel, and Du 2023). However, such methods can be vulnerable to scenarios where incorrect answers from malicious or poor-performing agents can influence the overall decision (Amayuelas et al. 2024).
Figure 1: Overview of our Multi-LLM framework. LLM-based agents that have explored the dynamic household are independently asked a binary “Yes/No” question and must answer based on their observations. Only one of the agents has the correct answer (“Yes”) while the other two agents are wrong due to either not exploring the relevant portions of the house enough or being malicious. The central network, however from supervised training, has learned to weigh the decision of the first agent more when given context about the item, room, and timestep.
An ideal EQA model would learn to identify incorrect agents and avoid being influenced by their responses. Furthermore, we would like to ideally avoid the communication costs incurred by debate-based approaches.
Main Results: To address the issues presented above, we apply a Multi-LLM system to solve this embodied task. We instruct the agents to answer based on their observations a set of binary embodied questions about the household. We then train a Central Answer Model (CAM) on the answers of independent agents with supervised learning. The network directly outputs the final answer to a given question and does not require any communication between agents. Furthermore, our model learns which agents to rely on while producing its final answer, which reduces the end user’s effort in having to individually figure out the best agent response. Our novel contributions include:
@@
  • Multi-LLM QA Framework with Embodied Navigation: We train a Central Answer Model (CAM) for Question Answering in an Embodied Multi-Agent setting. We depart from prior EQA setups where the agents explore after being asked a question. Rather we have the agents explore before being asked a set of question on their environment. CAM is a classifier that aggregates multiple agent responses to predict an answer. We generate queries for various Matterport 3D (Chang et al. 2017) environments. We train CAM on the labeled query dataset with multiple machine learning methods and show that our best model can achieve up to 46% higher accuracy than a majority vote and debate aggregation methods. While prior work has used a central answer module to output a final answer from a mixture of answers from multiple LLM experts (Puerto, Şahin, and Gurevych 2023; Si et al. 2023; Jiang, Ren, and Lin 2023), we are the first to apply the framework to embodied agents that must first explore an environment.
@@
  • Integration with Exploration Systems: We evaluate our multi-agent framework on data gathered using an LLM-based exploration method (Dorbala, Mullen Jr, and Manocha 2023) on multiple agents in Matterport3D, demonstrating our system can work in conjunction with SOTA exploration methods in unknown environments.
@@
  • Examining Agent and Feature Reliance: We present a feature importance analysis of CAM reliance on each independent agent using the metric, permutation feature importance (PFI) (Breiman 2001; Fisher, Rudin, and Dominici 2018). We also look at the agreement between the individual agent and the final answer of the central answer model and the baselines. Similarly to Si et al. (2023), some of our methods for the central answer model are interpretable and provide a Decision Tree visualization of one our models.
Reference Exploration Central Module Multi-Agent Feature Importance
(Majumdar et al. 2024)
(Puerto, Şahin, and Gurevych 2023)
Our Approach
Table 1: Table of our contributions and novelties compared to recent prior work. This work is the first to tackle Embodied EQA with LLM-based Multi-agent system and to provide a feature importance analysis on the framework.
Related Works
Embodied Question Answering and Exploration
@@
Embodied Question Answering (EQA) was first introduced by Das et al. (2018) and has become a popular task for robotics (Gordon et al. 2018; Gordon, Fox, and Farhadi 2019; Duan et al. 2022a; Sima et al. 2023; Wijmans et al. 2019; Chen et al. 2019; Luo et al. 2019), where traditionally an agent is given a query such as “was there a plant in the living room?’ and explores its environment, in this case a house, to answer the question. Zhu et al. (2023) had alternated between an exploration phase and then a question phase and repeated the process until the query was answered. Further work has extended EQA to scenarios where the final answer revolves around multiple objects (Yu et al. 2019), and where a target object is not explicitly stated in the question (Tan et al. 2023). Tan et al. (2020) introduced multi-agent exploration and cooperation for question-answering in interactive environments. Dorbala et al. (2024a) provided a dataset of subjective or situational queries and was the first to use a generative approach for query creation, and Majumdar et al. (2024) designed the first open-vocabulary EQA dataset to test and benchmark foundation models on this task. Recently, Singh, Pavlakos, and Stamoulis (2024) studied the zero-shot LLM performance on visual QA with indoor house scenes. However, the LLM is given the scenes without any exploration of the environment. We evaluate our Multi-LLM approach with a SOTA LLM-based exploration method for observation gathering introduced by Dorbala, Mullen Jr, and Manocha (2023).
@@
Note: In our work, we deviate from the traditional EQA approach. We first have an exploration gathering phase, then answer a set of questions on the same environment. We do so for two reasons: 1) to provide a setup to better compare against Multi-LLM QA baselines: majority voting and debating; 2) we believe this setup is better suited for the scenario where the user asks a question about the items in their homes and quickly needs an answer.
Multi-LLM Systems and Experts
@@
Research in Multi-LLM systems have grown in popularity with the rise of foundation models for areas such as task planning (Chen et al. 2023b), application development (Wu et al. 2023), and chatbots (Duan et al. 2023). Extensive studies have focused on their abilities to cooperate. Prior work has used multi-LLM debate to reach a consensus (Chan et al. 2023; Du et al. 2023; Guo et al. 2024a; Chen, Saha, and Bansal 2023). Chen et al. (2023a) analyzes the consensus-seeking debate between multiple LLM agents with different personalities, and Amayuelas et al. (2024) investigates the susceptibility of a multi-LLM discussion to an adversarial agent that purposefully tries to lead an incorrect answer choice. For embodied AI, Guo et al. (2024b) have shown that multiple LLM-based agents can cooperate in teams by designating a leader for increased task efficiency. Other work have looked into using voting metrics of independent LLM-agents to arrive at a final answer (Li et al. 2024; Wei et al. 2022; Wang et al. 2022). Our work aims to remove the need for communication by training a central classifier on independent answers from multiple agents. Training a classifier in a supervised manner helps our approach become robust to unreliable agents that may influence overall answers in voting schemes.
@@
A central architecture to output a final answer from multiple answers from different LLM experts has been explored previously (Puerto, Şahin, and Gurevych 2023; Si et al. 2023; Jiang, Ren, and Lin 2023). Our approach differs from these prior works as the various LLM agents we utilize become experts in the environments from an embodied exploration phase and we show it can be used with off-the-shelf exploration methods. We believe that our approach can be considered a variation of Mixture-of-Experts (MoE) (Eigen, Ranzato, and Sutskever 2013), as early MoE methods relied on training ML models like SVM to learn the weightings of different experts, as our CAM framework does. Our main difference is that the MoE module also takes into account answers from embodied agents explore the environment.
Figure 2: Overview of experimentation pipeline. [LEFT] We have an LLM answer a collection of embodied questions about the Matterport3D environment using the observations collected and stored by a specific agent during an exploration phase. The answers from the LLM are stored offline. [RIGHT] In the central network training phase, we select multiple sets of answers from different agents to train the classifier with the ground-truth answer with supervised learning.
Multi-LLM System with Central Answer Model (CAM) for EQA
Problem Formulation
@@
In this section, we provide a mathematical formulation of our specific EQA setting to characterize the EQA problem we are tackling in this work. We consider a tuple of $(Q,S^{K})$, where $Q$ represents the set of queries, where each sample is a vector $\left[o,r,y\right]$ with object $o$ and room $r$ representing the question “Is there an $o$ in the $r$?”, and $y\in\{0,1\}$ is the binary ground-truth answer to the question. $S$ is the set of $K$ embodied agents that independently answer the questions. Let $s_{k}^{(o,r)}\in\{0,1\}$ be the independent answer to query $(o,r)$ of the $k$-th agent. Let $f$ be a function such that $f\left(o,r,\cup_{k=1}^{K}s_{k}^{(o,r)}\right)=\hat{y}\in\{0,1\}$. Given the query context $(o,r)$ and the independent agent answers to the context, $f$ should produce a final binary prediction to whether or not object $o$ is in room $r$. In this work, we present CAM as a viable method of the function $f$ over baselines majority voting and debating. We now describe CAM in the next section.
Our Approach
@@
We implement a two-step procedure that comprises a question-answering phase, and a central model training phase. The question-answering phase is an agent-specific step while the classification-training step relies on combining the intermediate results of the question-answering phases from various agents. Before starting the question-answering phase, the agents explored the environment to gather observations. Later, we examine the performance of our approach when using a LLM-based SOTA method for exploration. We describe both phases below.
Question-Answering Phase
@@
In this phase, we generate a dataset of binary embodied questions specific to a Matterport 3D environment. Matterport3D is a large-scale dataset popular for scene-understanding tasks and embodied AI research, containing RGB-D panoramic views of various building environments like homes. The data is stored as a graph for each environment where the nodes are different physical locations within the home environment that an embodied agent can occupy. We use a topological graph environment because we are only concerned with high-level exploration policy in the house environment. Our approach can be used with other similar environments such as HM3D (Ramakrishnan et al. 2021).
@@
Central Model Training Phase
To give the final answer to each question, we train a model $f$ to give a final, binary “Yes/No” output. The model input dimension is $2+K$, where $K$ is the number of agents. The first two inputs account for the vectorized question in the form of “[object, room]”, $(o,r)$. The rest of the $K$ inputs come from entering the independent answers from each of the agents chosen for the multi-LLM system, $\cup_{k=1}^{K}s_{k}^{(o,r)}$. $f$ then outputs a predicted answer $\hat{y}$ and is then optimized with the ground truth answer $y$. During inference, we input unseen questions and agent answers into the network to output a test-time answer. Due to the modularity of our approach, for each set of independent responses, we train multiple machine learning models to produce multiple central answer models to compare against baseline aggregation approaches.
Experimentation
We now describe our experimentation in detail. With these experiments, we aim to 1) measure the accuracy of CAM for EQA and show its performance over other aggregation baselines (Results section); 2) and 3) provide a feature importance analysis to highlight the benefit of utilizing independent agent responses in our approach (Feature Importance Analysis Section) In all experiments we use GPT-4-turbo.
Query Generation
We extract the embodied question bank from the Matterport3D environments. For each scan, we extract an offline list of observed items detected in each image of the environment using the GLIP object detection algorithm (Li et al. 2022). The items space observed by GLIP can be found in the Appendix in Table 6.
For each item and room combination, we formulate the question using the template described in Section Question-Answering Phase. Because these questions were extracted based on the observed items in the scan, we know the ground truth answer is “Yes”.
To create questions with ground-truth answers “No”, for each query generated with a “Yes” ground-truth, we formulate a question with the same item but with a random room that is not within the subset of rooms that contains that specific item.
Note: Our setup is highly dependent on the object detector used. Error can thus be introduced by the object detection with GLIP and then the LLM answer generation. Some outputs of GLIP included toilets in the living room and a TV in the hallway. When the LLM agent has a (room, item) in its observation memory in the system prompt, it will usually say Yes if asked about said room and item, but when it does not have it in memory, the questions are harder to answer, leading to error in answers generated by the independent LLM agent. More error analysis on agent answers will be studied in further in the Feature Importance Analysis section.
Evaluation
We use a SOTA algorithm, LGX (Dorbala, Mullen Jr, and Manocha 2023) for the exploration phase.
Metric. To measure and compare the performances of our CAM approach against baseline aggregation approaches, we calculate the inference time accuracy, or correctness, of the model:
$\frac{1}{N}\sum_{i=1}^{N}\mathbf{1}_{y_{i}=\hat{y}_{i}},$ (1)
where $y_{i}$ is the ground truth label of the test question, and $\hat{y}_{i}$ is the final answer prediction from CAM. We produce CAM models with various machine learning algorithms: Neural Network (NN), Random Forest (RF), Decision Tree (DT), XGBoost (XG), RBF SVM (SVM), Linear SVM (SVM-L), Logistic Regression (LR).
The NN was implemented with Pytorch using 3 linear layers. The dimension were 16, 8, 1 in that order with a Sigmoid activation function at the end to ensure the output was in between 0 and 1. XGBoost used the default settings from the xgboost library package. RF from the sklearn library used $1000$ trees, and all other methods used the default settings from sklearn.
Baselines
We compare these different variations of CAM against two other forms of the aggregation methods used for ensemble LLM methods. Neither method is learning-based and requires no training set. We describe them below:

### v5

# Multi-LLM QA with Embodied Exploration

@@ Abstract
Large language models (LLMs) have grown in popularity due to their natural language interface and pre trained knowledge, leading to rapidly increasing success in question-answering (QA) tasks. More recently, multi-agent systems with LLM-based agents (Multi-LLM) have been utilized increasingly more for QA. In these scenarios, the models may each answer the question and reach a consensus or each model is specialized to answer different domain questions. However, most prior work dealing with Multi-LLM QA has focused on scenarios where the models are asked in a zero-shot manner or are given information sources to extract the answer. For question answering of an unknown environment, embodied exploration of the environment is first needed to answer the question. This skill is necessary for personalizing embodied AI to environments such as households. There is a lack of insight into whether a Multi-LLM system can handle question-answering based on observations from embodied exploration. In this work, we address this gap by investigating the use of Multi-Embodied LLM Explorers (MELE) for QA in an unknown environment. Multiple LLM-based agents independently explore and then answer queries about a household environment. We analyze different aggregation methods to generate a single, final answer for each query: debating, majority voting, and training a central answer module (CAM). Using CAM, we observe a $46\%$ higher accuracy compared against the other non-learning-based aggregation methods. We provide code and the query dataset for further research.
Multi-LLM QA with Embodied Exploration
Bhrij Patel University of Maryland, College Park                        Vishnu Sashank Dorbala University of Maryland, College Park Dinesh Manocha University of Maryland, College Park                        Amrit Singh Bedi University of Central Florida
1 Introduction
Large Language Models (LLMs) (Achiam et al., 2023) become more embedded into society, with one of the most popular tasks being question-answering (QA) (Rawal et al., 2024; Kamalloo et al., 2023). Further research has investigated utilizing an ensemble of LLM models (Multi-LLM) for the ability to discuss (Li et al., 2023; Chen et al., 2023b), answer questions from large data sources and route questions to models with specialized knowledge (Mixture of Experts) (Li et al., 2024).
However, in these prior works on ensemble LLMs for QA, the questions are asked in a zero-shot manner or information is provided to the models. For QA of a physical environment, an embodied agent must first explore an environment, such as a household, to answer questions based on its observations. It has gained considerable interest in the past few years due to its applications for in-home robots and personalized assistants Das et al. (2018); Gordon et al. (2018); Duan et al. (2022a); Sima et al. (2023); Luo et al. (2023); Dorbala et al. (2024a); Zhu et al. (2023). Possessing zero-shot common-sense reasoning of relational understandings of household items is crucial for LLMs to perform embodied exploration in unknown environments Dorbala et al. (2023, 2024b); Majumdar et al. (2024).
Figure 1: Overview of our MELE framework. LLM-based agents that have explored the dynamic household are independently asked a binary “Yes/No” question and must answer based on their observations. Only one of the agents has the correct answer (“No”) while the other two agents are wrong. The aggregator function then takes in these responses to decide a final response.
@@ 1 Introduction
In existing works, only a single agent explores the environment with a limited field of view, inducing low coverage and high exploration costs. In contrast, for an LLM-based embodied agent to succeed in EQA, efficient exploration of the environment is needed as the questions will center around the state of the environment Das et al. (2018); Yu et al. (2019); Zhu et al. (2023); Majumdar et al. (2024). The rise of Embodied Artificial Intelligence (EAI) Duan et al. (2022b) brings up the possibility of co-habited spaces with multiple agents and humans working in conjunction Puig et al. (2023).
@@
With this trend, we believe investigating the abilities of multiple embodied LLM agents to answer questions on their environment becomes increasingly crucial for the personalization of AI. We consider a similar co-inhabited setup and propose using a team comprising multiple embodied LLM-based explorers. An ensemble system of explorers allows for simultaneous exploration of the environment in the same amount of time.
@@
However, there are also challenges when introducing a multi-agent setup. A key challenge in a multi-agent setup is with conflicting responses when agents are asked the same query. Since LLMs are blackboxes capable of generating free-form natural language, their responses to the same query may significantly vary Gal and Grosz (2022); Wang et al. (2024b). In an embodied setting, such responses are further skewed by partial observations due to limited exploration. Prior work in ensemble LLM methods tackle this issue using weighted sum or voting schemes Li et al. (2024); Wei et al. (2022); Wang et al. (2022); Yang et al. (2024) and consensus-reaching debates between the LLM agents Chen et al. (2023a); Du et al. (2023); Xiong et al. (2023); Guo et al. (2024a); Chan et al. (2023); Li et al. (2023). However, such methods can be vulnerable to scenarios where incorrect answers from malicious or poor-performing agents can influence the overall decision Amayuelas et al. (2024).
Main Results: To investigate the use of ensemble LLM for QA with embodied exploration, we apply a system of Multi-Embodied LLM Explorers (MELE) to solve this embodied task. We instruct the agents to answer based on their observations a set of binary embodied questions about the household. We then use various aggregation methods to output a final answer based on the independent responses from the agents: majority voting, debating, and training a supervised central answer model (CAM). Majority voting and debating are common aggregation methods in existing literature and we propose investigating training CAM for more robustness to adversarial or incorrect agents. Our novel contributions include:
@@
  • Multi-LLM QA Framework with Embodied Navigation: We use multiple independent embodied LLM agents that explore the environment individually generate a response and use various aggregator functions to predict a final response.  [^1] Link to anonymous repository: https://anonymous.4open.science/r/CAM-embodied-llm-54CE/ We depart from prior Embodied Question-Answering (EQA) Das et al. (2018) setups where the agents explore after being asked a question. Rather we have the agents explore before being asked a set of questions on their environment. We generate queries for various Matterport 3D Chang et al. (2017) environments. We train CAM on the labeled query dataset with multiple machine learning methods and show that the best CAM-based method can achieve up to 46% higher accuracy than a majority vote and debate aggregation methods. While prior work has used a central answer module to output a final answer from a mixture of answers from multiple LLM experts Puerto et al. (2023); Si et al. (2023); Jiang et al. (2023), we are the first to apply the framework to embodied agents that must first explore an environment.
@@
  • Integration with Exploration Systems: We evaluate our multi-agent framework on data gathered using an LLM-based exploration method Dorbala et al. (2023) on multiple agents in Matterport3D, demonstrating our system can work in conjunction with SOTA exploration methods in unknown environments.
@@
  • Examining Agent and Feature Reliance: We present a feature importance analysis of CAM reliance on each independent agent using the metric, permutation feature importance (PFI) Breiman (2001); Fisher et al. (2018). We also look at the agreement between the individual agent and the final answer of the central answer model and the baselines. Similarly to Si et al. (2023), some of our methods for the central answer model are interpretable and provide a Decision Tree visualization of one of our models.
2 Related Works
2.1 Embodied Question Answering and Exploration
@@
Embodied Question Answering (EQA) was first introduced by Das et al. (2018) and has become a popular task for robotics (Gordon et al., 2018, 2019; Duan et al., 2022a; Sima et al., 2023; Wijmans et al., 2019; Chen et al., 2019; Luo et al., 2019), where traditionally an agent is given a query such as “was there a plant in the living room?’ and explores its environment, in this case a house, to answer the question. Zhu et al. (2023) had alternated between an exploration phase and then a question phase and repeated the process until the query was answered. Further work has extended EQA to scenarios where the final answer revolves around multiple objects Yu et al. (2019), and where a target object is not explicitly stated in the question Tan et al. (2023). Tan et al. (2020) introduced multi-agent exploration and cooperation for question-answering in interactive environments. Dorbala et al. (2024a) provided a dataset of subjective or situational queries and was the first to use a generative approach for query creation, and Majumdar et al. (2024) designed the first open-vocabulary EQA dataset to test and benchmark foundation models on this task. Recently, Singh et al. (2024) studied the zero-shot LLM performance on visual QA with indoor house scenes. However, the LLM is given the scenes without any exploration of the environment. We evaluate our Multi-LLM approach with a SOTA LLM-based exploration method for observation gathering introduced by Dorbala et al. (2023).
@@
Note: In our work, we deviate from the traditional EQA approach. We first have an exploration gathering phase, then answer a set of questions in the same environment. We do so for two reasons: 1) to provide a setup that better resembles prior work in LLM-based MoE QA (Li et al., 2024) 2) we believe this setup is better suited for the scenario where the user asks a question about the items in their homes and quickly needs an answer.
2.2 Systems of Multiple LLM-based Agents
@@
Research in LLM-based systems have grown in popularity with the rise of foundation models for areas such as task planning Chen et al. (2023c), application development Wu et al. (2023), and chatbots Duan et al. (2023). Extensive studies have focused on their abilities to cooperate. Prior work has used debate between LLM models to reach a consensus (Chan et al., 2023; Du et al., 2023; Guo et al., 2024a; Chen et al., 2023b). Chen et al. (2023a) analyzes the consensus-seeking debate between multiple LLM agents with different personalities, and Amayuelas et al. (2024) investigates the susceptibility of a LLM-based discussion to an adversarial agent that purposefully tries to lead an incorrect answer choice. For embodied AI, Guo et al. (2024b) have shown that multiple LLM-based agents can cooperate in teams by designating a leader for increased task efficiency. Other work have looked into using voting metrics of independent LLM-agents to arrive at a final answer (Li et al., 2024; Wei et al., 2022; Wang et al., 2022). Our work aims to remove the need for communication by training a central classifier on independent answers from multiple agents. Training a classifier in a supervised manner helps our approach become robust to unreliable agents that may influence overall answers in voting schemes.
@@
A central architecture to output a final answer from multiple answers from different LLM experts has been explored previously Puerto et al. (2023); Si et al. (2023); Jiang et al. (2023). Our approach differs from these prior works as the various LLM agents we utilize become experts in the environments from an embodied exploration phase and we show it can be used with off-the-shelf exploration methods. Furthermore, Wang et al. (2024a) introduced the idea of a Mixture of Agents to enhance task performance of LLMs. We specifically look at a mixture of embodied agents that explore a unknown environment.
3 Multi-Embodied LLM Explorers
3.1 Problem Formulation
@@
In this section, we provide a mathematical formulation of our specific EQA setting to characterize the EQA problem we are tackling in this work. We consider a tuple of $(Q,S^{K})$, where $Q$ represents the set of queries, where each sample is a vector $\left[o,r,y\right]$ with object $o$ and room $r$ representing the question “Is there an $o$ in the $r$?”, and $y\in\{0,1\}$ is the binary ground-truth answer to the question. $S$ is the set of $K$ embodied agents that independently answer the questions. Let $s_{k}^{(o,r)}\in\{0,1\}$ be the independent answer to query $(o,r)$ of the $k$-th agent. Let $f$ be a aggregation function such that $f\left(o,r,\cup_{k=1}^{K}s_{k}^{(o,r)}\right)=\hat{y}\in\{0,1\}$. Given the query context $(o,r)$ and the independent agent answers to the context, $f$ should produce a final binary prediction of whether object $o$ is in room $r$.
3.2 Our Approach
@@
We implement MELE as a two-step procedure that comprises a question-answering phase, and a aggregation phase. The question-answering phase is an agent-specific step while the classification-training step relies on combining the intermediate results of the question-answering phases from various agents. Before starting the question-answering phase, the agents explored the environment to gather observations. Later, we examine the performance of MELE when using a LLM-based SOTA method for exploration.
3.3 Question-Answering Phase
@@
In this phase, we generate a dataset of binary embodied questions specific to a Matterport 3D environment. Matterport3D is a large-scale dataset popular for scene-understanding tasks and embodied AI research, containing RGB-D panoramic views of various building environments like homes. The data is stored as a graph for each environment where the nodes are different physical locations within the home environment that an embodied agent can occupy. We use a topological graph environment because we are only concerned with high-level exploration policy in the house environment. Our approach can be used with other similar environments such as HM3D Ramakrishnan et al. (2021).
@@
3.4 Aggregation Phase
We now describe the different methods used for aggregation function $f$.
3.4.1 Non-learning-based Aggregation
We first describe the non-learning based methods used commonly in ensemble LLM literature.